solve the system of equations by graphing
then classify the system
2x-y=5
2x+3y=-7
subtract line 2 from line 1
2x-y=5
-2x -3y=7
add both
0x-4y=12
y=-4
substitute this value in eq 1 or 2x--4=5
or 2x =1
or x=1/2
ans x=1/2 y=-4
the solution is _________________
is the system consistant or inconsistant?
is the system dependent or independent?

Answer by shree840(260)

2x-y=5
2x+3y=-7
subtract line 2 from line 1
2x-y=5
-2x -3y=7
add both
0x-4y=12
y=-4
substitute this value in eq 1 or 2x--4=5
or 2x =1
or x=1/2